# Can Community Notes Replace Professional Fact-Checkers?

### v1

@@ Abstract
Two commonly-employed strategies to combat the rise of misinformation on social media are (i) fact-checking by professional organisations and (ii) community moderation by platform users. Policy changes by Twitter/X and, more recently, Meta, signal a shift away from partnerships with fact-checking organisations and towards an increased reliance on crowdsourced community notes. However, the extent and nature of dependencies between fact-checking and helpful community notes remain unclear. To address these questions, we use language models to annotate a large corpus of Twitter/X community notes with attributes such as topic, cited sources, and whether they refute claims tied to broader misinformation narratives. Our analysis reveals that community notes cite fact-checking sources up to five times more than previously reported. Fact-checking is especially crucial for notes on posts linked to broader narratives, which are twice as likely to reference fact-checking sources compared to other sources. In conclusion, our results show that successful community moderation heavily relies on professional fact-checking.
@@ 1 Introduction
The proliferation of misinformation on social media (Arnold, 2020; Diakopoulos, 2020), along with the rise of generative AI (Augenstein et al., 2024) have led to increasing concerns about its current and future potential harms, (e.g., to health (Clemente et al., 2022)) and threats to democracy and political stability (Reglitz, 2022).
Fact-checkers play a crucial role in combatting misinformation (Graves, 2017), and in recent years, have partnered with social media platforms, e.g., Meta, YouTube, and TikTok, to tackle its spread on these platforms. However, due to the scale of misleading content shared online, community moderation (e.g., options to flag potential misinformation, group/server moderators) is often employed in parallel (Morrow et al., 2022), as a complementary approach (e.g., (Google, 2025); see also the practice of snoping (Pilarski et al., 2024)).
@@
The expansion of fact-checking projects in the last decade (Lauer and Graves, 2024), alongside their broader initiatives to curb misinformation (e.g., citizen media literacy programmes (Juneja and Mitra, 2022)) have been aided by partnerships with social media platforms such as Meta and Google (Graves and Anderson, 2020), which fund independent fact-checking agencies to fact-check potentially false claims on their platform. [^1] Fact-checkers provide a judgment of claim veracity and exert no influence on the platforms’ content moderation policies (Catalanello and Sanders, 2025).
However, political pressure and accusations of bias and censorship, and most recently, Meta’s announcement of its plans to end its partnerships with fact-checkers in the U.S. and implement a community moderation model (Meta, 2025), threatens the financial stability of fact-checking organisations, and hence, their ability to keep up with the increasing volume and sophistication of misinformation spread (Stencel et al., 2024; IFCN, 2024).
Meta’s recent policy shift also implies that these two strategies (fact-checking and community notes) are independent and in opposition, rather than two complementary strategies of tackling online misinformation. In this paper, we examine Twitter/X community notes as a case study to understand how fact-checking is used in community notes. Specifically, we investigate the following two questions: (RQ1) To what extent do community notes rely on the work of professional fact-checkers? and (RQ2) What are the traits of posts and notes that rely on fact-checking sources? Studying the relationship between fact-checking and community notes is vital for understanding the shared role of expert and community-driven fact-checking in the global information ecosystem.
@@ 2.1.1 Characteristics of Community Notes
Sources in notes. Analyses have indicated that notes were rated more helpful if they link to ‘trustworthy’ sources (Pröllochs, 2022) and that the majority of sources cited by notes were ‘trustworthy’ left-leaning news outlets. A recent study finds that 55% of URLs used in notes were related to news websites, 18% to research, 9% to social media, 9% to encyclopedic sources, but just 1.2% to fact-checking sources (Kangur et al., 2024).
@@
Limitations of notes. Only 11% of submitted notes reach ‘helpful’ status (i.e., shown to users) by achieving a cross-perspective (Renault et al., 2024; Wirtschafter and Majumder, 2023), and the time frame for notes to reach the algorithm’s required agreement level (15.5 hours on average) limits its capacity to halt misinformation spread (Renault et al., 2024). Additional concerns about the notes’ efficacy highlight their indifference to the expertise needed for certain claims and reliance on subjective helpfulness rather than objective facts, free labour and inadequate support and guardrails regarding explicit content (Gilbert, 2025).
@@
Our work provides novel insights into the targets, sources and limitations of community notes by shedding light on the relationship between notes and professional fact-checking. Namely, we study the extent to which fact-checking sources form the basis of note-writers’ efforts to counter misinformation and identify the strategies they employ.
2.1.2 Impact of Community Notes on misinformation spread
People shown community notes alongside misleading social media posts were more accurate in identifying misleading posts, and the notes were judged to be more trustworthy than context-free misinformation flags (e.g., "Checked by fact-checkers" or "Checked by other social media users"), regardless of (US-centric) political beliefs Drolsbach et al. (2024). People shown either community notes or related news article suggestions were both less likely to to believe and report misleading information compared to a control group: community notes were more effective in reducing belief and sharing intention for positive rumours, while articles were more effective for negative rumours Kankham and Hou (2024). On the other hand, displaying community notes leads users to post more negative and angry replies to misleading posts Chuai et al. (2024b), while crowd workers are also prone to cognitive biases, such as overestimating a statement’s truthfulness the more they liked its claimant, and general overconfidence in their ability to ascertain truthful statements Draws et al. (2022).
2.2 Professional fact-checking and community note practices
Although fact-checks and community notes share similarities in how they address misleading claims, they also differ in key elements of practice and techniques of persuasion and communication (Kankham and Hou, 2024). Fact-checking typically involves the analysis and verification of public claims (e.g., statements in news reports and social media). In addition to verifying claims, in recent years many fact-checking organisations have also assumed a wider role in combatting misinformation spread, conducting long-term investigative journalism projects and citizen media literacy programs (Juneja and Mitra, 2022). Professional fact-checkers in organisations signatory to the International Fact-Checking Network (IFCN) follow a rigorous set of principles and transparency commitments. [^2] https://www.ifcncodeofprinciples.poynter.org/the-commitments In contrast, any platform user can contribute to community notes under anonymity, and the rating approach relies on the ’wisdom of crowds’, with little oversight or transparency regarding the biases of the note-writers. Numerous studies have documented the structured workflow that fact-checkers follow: (i) claim selection; (ii) collecting evidence; (iii) deciding on a verdict; and (iv) writing the fact-checking article (Graves, 2017; Micallef et al., 2022; Warren et al., 2025). Fact-checking articles, which are subject to multiple rounds of editorial scrutiny, are more formal and standardised in tone and style than community notes, which vary considerably. Fact-checkers must rely on credible sources and evidence to convince the reader, while community note writers may employ a range of persuasion techniques, such as appeals to emotion or other logical fallacies. Moreover, community notes typically serve as direct rebuttals to misleading posts, while fact-checking articles may address a more general claim than is expressed in a specific post. Finally, fact-checking articles are a one-way exchange, while community notes represent a more horizontal and interactive dialogue between writer and recipient of the fact-check (Kankham and Hou, 2024).
Our work builds on current understanding of the relationship between professional fact-checking and amateur community moderation by examining the extent to which community note writers deploy the work of professional fact-checkers in their notes.
@@ 3 Dataset
Figure 2: The categories of links used by Community notes’ authors as a source.
We download files containing all community notes and their metadata from the official website, [^3] https://communitynotes.x.com/guide/en/under-the-hood/download-data which amounts to 1.5M notes authored between January 28th 2021 and January 6th 2025. Of these, a total of 135K are rated by the community as ‘Helpful’, 51K are rated ‘Not helpful’, and 1.3M are unpublished, i.e., did not receive enough community ratings to reach a verdict. See Fig. 6 in App. A for statistics.
We filter the notes as follows. First, we remove 526K non-English notes, which we identify by applying the language detection library fast-langdetect. [^4] https://github.com/LlmKira/fast-langdetect Then, we further filter 268K “unnecessary” notes—notes attached to tweets that are classified by the community as “not misleading”. Finally, to focus only on notes that are used to address misinformation, we filter out 44K notes that contain one of the words “ad”, “spam”, or “phishing”. Following these filtration steps, we are left with a dataset containing 664K notes.
The next step involves categorising the sources that the note authors use to support their claims. First, we use regex to extract all the URLs found in the notes. See Tab. 4 in App. A for a list of the top-100 most common domains. We classify each URL in our dataset of 664K notes into one of 13 categories (detailed in Fig. 2) using the pipeline described below.
@@
    Check whether the domain name of the URL is found in a manually curated list of domains of professional fact-checking organisations (See Tab. 3 in App. A for the full list). If so, classify the URL as “fact-checking”.
@@
    Otherwise, search for paraphrases of the word “fact-check” in the URL, [^5] These URLs mostly link to the fact-checking divisions of news outlets, e.g., https://apnews.com/article/fact-checking-909101991741 and classify it as “fact-checking” if a match was found.
@@
    Otherwise, use GPT-4666Version gpt-4o-2024-08-06. to classify the domain name into one of the 13 categories. LABEL:lst:prompt_link in App. B details the prompt we used.
@@
    Finally, if GPT-4 fails or outputs an unknown category, label the URL as “unknown”.
@@
Moreover, we further subsample the notes for performing the in-depth analysis required for answering RQ2 (§ 4.2). From the notes rated as ‘Helpful’, we sample 3.5K notes with a “Fact-checking” source and a random sample of 22K additional notes. We then used web crawling to scrape the text of the posts to which these notes were attached. We name this subset $\mathcal{S}_{\text{text}}$ for simplicity.
@@ 4.1 RQ1: To what degree do community notes rely on fact-checkers?
According to Fig. 2, at least 5% of all English community notes contain an external link to professional fact-checkers. This number grows to 7% when only considering notes rated as ‘helpful’ (Fig. 7 in App. A). Conversely, only 1% of notes rated as ‘not helpful’ contain a fact-checking source (Fig. 8 in App. A). These figures are significantly larger than what was reported in previous studies (1.2% (Kangur et al., 2024)), possibly because Kangur et al. (2024) utilise a smaller dataset of fact-checking agencies and classify fact-checking divisions of popular journals as “news”. The results imply that notes incorporating fact-checking sources are generally considered more helpful.
We further assess whether notes with fact-checking sources are perceived to be of higher quality by analysing individual user ratings of notes both with and without such sources. Specifically, we collect user ratings for a balanced (i.e., including of a fact-checking source or not) sample of 20K notes rated by at least 50 users, and calculated the average ratings for the notes. As can be seen in Fig. 9 in App. A, community notes with fact-checking sources are generally rated higher than their counterparts. Interestingly, while notes with fact-checking links are more likely to be regarded as having a good source (higher HelpfulGoodSources), they are also more likely to be rated as notHelpfulSourcesMissingOrUnreliable. Tab. 5 in App. A contains a sample of such notes.
@@ 4.2 RQ2: What are the traits of posts and notes that rely on fact-checking sources?
Tweet Note
misleadingUnverifiedClaimAsFact
misleadingOutdatedInformation
misleadingFactualError
misleadingSatire
Fact Checking source
@cnviolations I swear community notes are the only good thing Elon added since he bought Twitter. Community notes was first launched under former Twitter CEO Jack Dorsey in 2021 under the name of “Birdwatch”. The only thing Elon Musk did was that he renamed the feature to community notes. https://blog.twitter.com/en_us/topics/product/2021/introducing-birdwatch-a-community-based-approach-to-misinformation https://www.reuters.com/article/factcheck-elon-birdwatch-idUSL1N31Z2VG/
Thailand will become the first country to make the contract null and void, meaning that Pfizer will become responsible for all vaccine injuries … Thailand has no plans to void its Pfizer COVID vaccine contract, an official with the country’s National Vaccine Institute said. Thailand’s Department of Disease Control also rejected the claims as “fake news.” … https://apnews.com/article/fact-check-covid-vaccine-pfizer-thailand-203948163859
@@
We begin by performing a topic analysis, comparing topics of posts whose notes reference fact-checking sources to those citing other sources. To this end, we apply a strong zero-shot text classification model [^7] https://huggingface.co/r-f/ModernBERT-large-zeroshot-v1 with default settings. to our $\mathcal{S}_{\text{text}}$ subset by classifying spans of the form “Tweet:<POST TEXT>; Note <NOTE TEXT>” into one of 13 classes. The authors manually evaluated the quality of the classification results and considered it satisfactory. Notably (Fig. 5), fact-checking sources are more likely to be included in posts related to high-stakes issues such as health, science, and scams and less likely to be included in posts on tech or sports.
We then analyse annotations (binary attributes explaining the warrant for the note) by community note authors. Fig. 3 contains the full breakdown of annotations for notes with and without fact-checking sources. Notes containing a link to fact-checking sources are overrepresented in posts where unverified information is presented as a fact or when the post contains a factual error. Conversely, they are under-represented in posts with outdated information or satirical content. Tab. 1 contains a sample of such notes.
[…]
@@
[…]

### v2

@@ Abstract
Two commonly employed strategies to combat the rise of misinformation on social media are (i) fact-checking by professional organisations and (ii) community moderation by platform users. Policy changes by Twitter/X and, more recently, Meta, signal a shift away from partnerships with fact-checking organisations and towards an increased reliance on crowdsourced community notes. However, the extent and nature of dependencies between fact-checking and helpful community notes remain unclear. To address these questions, we use language models to annotate a large corpus of Twitter/X community notes with attributes such as topic, cited sources, and whether they refute claims tied to broader misinformation narratives. Our analysis reveals that community notes cite fact-checking sources up to five times more than previously reported. Fact-checking is especially crucial for notes on posts linked to broader narratives, which are twice as likely to reference fact-checking sources compared to other sources. Our results show that successful community moderation relies on professional fact-checking and highlight how citizen and professional fact-checking are deeply intertwined.
{NoHyper}* Equal contribution.
@@ 1 Introduction
The proliferation of misinformation on social media (Arnold, 2020; Diakopoulos, 2020), along with the rise of generative AI (Augenstein et al., 2024) have led to increasing concerns about people’s ability to access trustworthy and credible information, leading to potential harms to public health (Clemente et al., 2022), democracy, and political stability (Reglitz, 2022). Fact-checkers play a crucial role in combatting misinformation (Graves, 2017), and in recent years, have partnered with social media platforms, e.g., Meta, YouTube, and TikTok, to tackle its spread on these platforms. However, due to the scale of misleading content shared online, community moderation (e.g., options to flag potential misinformation, group/server moderators) is often employed in parallel (Morrow et al., 2022), as a complementary approach (e.g., (Google, 2025; TikTok, 2025); see also the practice of snoping (Pilarski et al., 2024)). The expansion of fact-checking projects in the last decade (Lauer and Graves, 2024), alongside their broader initiatives to curb misinformation (e.g., citizen media literacy programmes (Juneja and Mitra, 2022)) have been aided by partnerships with social media platforms such as Meta and Google (Graves and Anderson, 2020), which fund independent fact-checking agencies to fact-check potentially false claims on their platform. [^1] Fact-checkers provide a judgment of claim veracity and exert no influence on the platforms’ content moderation policies (Catalanello and Sanders, 2025). However, political pressure and accusations of bias and censorship, and most recently, Meta’s announcement of its plans to end its partnerships with fact-checkers in the U.S. and implement a community moderation model (Meta, 2025), threatens the financial stability of fact-checking organisations, and hence, their ability to keep up with the increasing volume and sophistication of misinformation spread (Stencel et al., 2024; IFCN, 2024).
Meta’s recent policy shift also implies that these two strategies (fact-checking and community notes) are independent and in opposition, rather than two complementary strategies of tackling online misinformation. In this paper, we examine Twitter/X community notes as a case study to understand how fact-checking is used in community notes. Specifically, we investigate the following two questions: (RQ1) To what extent do community notes rely on the work of professional fact-checkers? and (RQ2) What are the traits of posts and notes that rely on fact-checking sources? Studying the relationship between fact-checking and community notes is vital for understanding the shared role of expert and citizen-driven fact-checking in the global information ecosystem.
@@ 2 Background
Due to space constraints, only the most relevant related work is provided here. Additional relevant studies and background can be found in App. A.
@@ 2.1.1 Characteristics of Community Notes
Sources in notes. Analyses have showed that notes were rated more helpful if they link to ‘trustworthy’ sources and that the majority of sources cited by notes were ‘trustworthy’ left-leaning news outlets (Pröllochs, 2022). A recent study finds that 55% of URLs used in notes were related to news websites, 18% to research, 9% to social media, 9% to encyclopedic sources, but just 1% to fact-checking sources (Kangur et al., 2024).
@@
Limitations of notes. Only 11% of submitted notes reach ‘helpful’ status (i.e., shown to users) by achieving a cross-perspective (Renault et al., 2024; Wirtschafter and Majumder, 2023), and the time frame for notes to reach the algorithm’s required agreement level (15.5 hours on average) limits its capacity to halt misinformation spread (Renault et al., 2024). Posts related to partisan issues are particularly affected by these challenges (Allen et al., 2022). Additional concerns about the notes’ efficacy highlight their indifference to the expertise needed for certain claims and reliance on subjective helpfulness rather than objective facts, free labour and inadequate support and guardrails regarding explicit content (Gilbert, 2025).
@@
Our work provides novel insights into the targets, sources and limitations of community notes by shedding light on the relationship between notes and professional fact-checking. Closest to our work is a recent analysis of community notes written in 2024 by the fact-checking organisation Maldita (2025), who also studied the reliance of community notes on professional fact-checkers. They discover that fact-checking organisations are widely used as a reference by notes’ authors. The current work provides a more fine-grained analysis by studying the extent to which fact-checking sources form the basis of note-writers’ efforts to counter misinformation and identifying the strategies they employ.
@@ 3 Dataset
Figure 2: The categories of links used by Community notes’ authors as a source. a) all community notes; b) Community notes rated as ‘helpful’; c) community notes rated as ‘unhelpful’. Notice the ‘fact-checking’ category.
We download files containing all community notes and their metadata from the official website, [^2] https://communitynotes.x.com/guide/en/under-the-hood/download-data which amounts to 1.5M notes authored between January 28th 2021 and January 6th 2025. Of these, a total of 135K are rated by the community as ‘Helpful’, 51K are rated ‘Not helpful’, and 1.3M are unpublished, i.e., did not receive enough community ratings to reach a verdict. See Fig. 6 in App. B for statistics.
We filter the notes as follows. First, we remove 526K non-English notes, which we identify by applying the language detection library fast-langdetect. [^3] https://github.com/LlmKira/fast-langdetect Then, we further filter 268K ‘unnecessary’ notes—notes attached to tweets that are classified by the community as ‘not misleading’. Finally, to focus only on notes that are used to address misinformation, we filter out 44K notes that contain one of the words ‘ad’, ‘spam’, or ‘phishing’. Following these filtration steps, we are left with a dataset containing 664K notes.
The next step involves categorising the sources that the note authors use to support their claims. First, we use regex to extract all the URLs found in the notes. Importantly, a single note can include multiple external URLs as evidence. See Tab. 2 in App. B for a list of the top-100 most common domains. We classify each URL in our dataset of 664K notes into one of 13 categories (detailed in Fig. 2) using the pipeline described below.
@@
    Check whether the domain name of the URL is found in a manually curated list of domains of professional fact-checking organisations (See Tab. 3 in App. B for the full list). If so, classify the URL as ‘fact-checking’.
@@
    Otherwise, search for paraphrases of the word ‘fact-check’ in the URL, [^4] These URLs mostly link to the fact-checking divisions of news outlets, e.g., https://apnews.com/article/fact-checking-909101991741. and classify it as ‘fact-checking’ if a match was found.
@@
    Otherwise, use GPT-4555Version gpt-4o-2024-08-06. to classify the domain name into one of the 13 categories. LABEL:lst:prompt_link in App. C details the prompt we used.
@@
    Finally, if GPT-4 fails or outputs an unknown category, label the URL as ‘unknown’.
@@
Moreover, we further subsample the notes for performing the in-depth analysis required for answering RQ2 (§ 4.2). From the notes rated as ‘helpful’, we sample 3.5K notes with a ‘fact-checking’ source and a random sample of 22K additional notes. We then used web crawling to scrape the text of the posts to which these notes were attached. We name this subset $\mathcal{S}_{\text{text}}$ for simplicity.
@@ 4.1 RQ1: To what degree do community notes rely on fact-checkers?
According to Fig. 2.a at least 5% of all English community notes contain an external link to professional fact-checkers. This number grows to 7% when only considering notes rated as ‘helpful’ (Fig. 2.b). Conversely, only 1% of notes rated as ‘not helpful’ contain a fact-checking source (Fig. 2.c). These figures are significantly larger than what was reported in some previous studies (1.2% (Kangur et al., 2024)), possibly because Kangur et al. (2024) utilise a smaller dataset of fact-checking agencies and classify fact-checking divisions of popular journals as ‘news’. The results imply that notes incorporating fact-checking sources are generally considered more helpful.
We further assess whether notes with fact-checking sources are perceived to be of higher quality by analysing individual user ratings of notes both with and without such sources. Specifically, we collect user ratings for a balanced (i.e., including of a fact-checking source or not) sample of 20K notes rated by at least 50 users, and calculated the average ratings for the notes. As can be seen in Fig. 7 in App. B, community notes with fact-checking sources are generally rated higher than their counterparts. Interestingly, while notes with fact-checking links are more likely to be regarded as having a good source (higher HelpfulGoodSources), they are also more likely to be rated as notHelpfulSourcesMissingOrUnreliable. Tab. 5 in App. B contains a sample of such notes.
@@ 4.2 RQ2: What are the traits of posts and notes that rely on fact-checking sources?
We begin by performing a topic analysis, comparing topics of posts whose notes reference fact-checking sources to those citing other sources. To this end, we apply a strong zero-shot text classification model [^6] https://huggingface.co/r-f/ModernBERT-large-zeroshot-v1 with default settings. to our $\mathcal{S}_{\text{text}}$ subset by classifying spans of the form “Tweet:<POST TEXT>; Note <NOTE TEXT>” into one of 13 classes. The authors manually evaluated the quality of the classification results and considered it satisfactory, with the model predicting the correct class in $90\%$ of the cases. Most of the incorrect predictions involved the ‘technology’ category, with sentences such as “this is a fake image that was created with AI” incorrectly labelled as ‘technology’. Notably (Fig. 5), fact-checking sources are more likely to be included in posts related to high-stakes issues such as health, science, and scams and less likely to be included in posts on tech or sports.
We then analyse annotations (binary attributes explaining the warrant for the note) by community note authors. Fig. 3 contains the full breakdown of annotations for notes with and without fact-checking sources. Notes containing a link to fact-checking sources are overrepresented in posts where unverified information is presented as a fact or when the post contains a factual error. Conversely, they are under-represented in posts with outdated information or satirical content. Tab. 4 in App. B contains a sample of such notes.
These results indicate that community note-writers adapt their strategies based on the stakes and scope of the claim, and the depth of research needed to counter misinformation. We hypothesise that they are more likely to rely on external fact-checking when refuting complex or unverifiable claims (Wuehrl et al., 2024), as well as claims related to conspiracy theories [^7] For example, the claim “Michelle Obama is a male”. or broader narratives [^8] These are tweets that perpetuate broader misinformation narratives but do not necessarily contain an explicit conspiracy theory, e.g., the false claim “most immigrants remain firmly dependent on Welfare”. which cannot be fully addressed in the scope of a note. Conversely, claims involving misleading media can often be debunked with examples alone, making fact-checking sources unnecessary. To investigate this hypothesis, the authors of this paper manually annotated 400 <post, note> pairs from $\mathcal{S}_{\text{text}}$ with attributes related to the complexity of the claims and how community notes address them. (see § C.1 for annotation guidelines). The results (Fig. 4.a) support our hypothesis. Claims related to broader narratives or conspiracy theories [^9] We condense broader narratives and conspiracy theories together for simplicity. Similar trends can be seen when they are analysed separately. are much more likely to include a link to a fact-checking source. In contrast, other types of claims are more likely to be addressed by providing missing context or by invalidating the credibility of the claim’s source. Additionally, Fig. 4.b depicts the different ways in which fact-checking sources are used to debunk claims. It demonstrates how such sources are rarely used to provide missing context but rather focus on discrediting sources of claims and providing scientific evidence.
@@
We extend the manual annotation to an LLM-based analysis of 8K balanced <post, note> pairs from $\mathcal{S}_{\text{text}}$. We task OpenAI’s GPT-4101010Version gpt-4o-2024-08-06. with determining whether a pair relates to a broader narrative or a conspiracy theory. LABEL:lst:prompt_conspiracy in App. C details the prompt we used. To evaluate model accuracy, two authors independently labelled 100 balanced pairs, achieving an agreement rate of $0.88$ and resolving disagreements through discussion. The model attained an $F_{1}$ score of $0.85$—strong performance for this challenging task. The results (Tab. 1) support our hypothesis: <post, note> pairs related to a broader narrative or conspiracy theory are twice as likely to cite fact-checking sources compared to other sources. In contrast, other pairs are nearly 30% less likely to do so. These findings also highlight the prevalence of such claims and further underscore the importance of fact-checking in combating complex misinformation narratives.